# MASA: Rethinking the Representational Bottleneck in LoRA with Multi-A Shared Adaptation

### v1

@@ Introduction
As discussed above, the single matrix $A$ in standard LoRA and its asymmetric variants (e.g., HydraLoRA) may become a representational bottleneck, limiting the model’s capacity to extract diverse and nuanced features. Our strategy is to reallocate the extra capacity to the feature extraction stage. The Multi-$A$ Expert (MAE) block replaces the single matrix $A$ with an ensemble of $N$ expert matrices, $\{A_{1},A_{2},\dots,A_{N}\}$, where each $A_{i}\in\mathbb{R}^{r\times d_{in}}$. These experts operate in parallel, enabling the model to capture more diverse and fine-grained features from the input. We hypothesize that each expert matrix learns to specialize in distinct semantic subspaces. This functional specialization is empirically supported by t-SNE (Maaten and Hinton 2008) visualizations in Fig. 2, which illustrate that different $A_{i}$ matrices map the same inputs to separable feature clusters. The outputs of these experts are aggregated and then projected back to the original dimension by a single restoration matrix $B\in\mathbb{R}^{d_{out}\times r}$. We aggregate expert features via summation, a simple yet effective approach that maintains parameter efficiency and training stability (Xu et al. ), with detailed rationale provided in Appendix B.
@@
Asymmetric Cross-layer Sharing
To mitigate the potential parameter increase from using multiple $A$-matrices and to reduce their redundancy across transformer layers, we introduce an asymmetric cross-layer sharing mechanism for efficient adaptation. Unlike other sharing methods (VB-LoRA (Li, Han, and Ji 2024), BSLoRA (Zhou et al. 2025)) that typically share the symmetric LoRA module (both $A$ and $B$), our approach introduces an asymmetric sharing strategy: we only share the ensemble of $A$-matrices across adjacent layers, while allowing each layer to retain its own independent, trainable $B$-matrix.
To substantiate our design, we conducted an empirical analysis of inter-layer activation similarity using Centered Kernel Alignment (CKA) (Kornblith et al. 2019), a robust method for comparing neural network representations. As presented in Fig. 3, we applied CKA to measure the similarity between activation values from adjacent layers of a fine-tuned model, specifically comparing the $A$-matrix outputs ($Ax$) against the final LoRA increment activations ($BAx$). The analysis reveals a distinct pattern: The $A$-matrix outputs maintain a significantly higher similarity across layers than the full increments. This observation is quantified by the aggregated plots, which confirm a consistently higher average similarity score for the $A$-matrix outputs across all attention projections ($q,k,v$). This evidence indicates that 1) $A$-matrices learn to produce generalizable cross-layer representations, making them ideal candidates for sharing; and 2) $B$-matrices, conversely, perform more layer-specific transformations, which are critical for preserving the unique functionality of each layer.
@@
The Whole MASA Framework
To compromise performance and efficiency, we combine the Multi-$A$ Expert block with our asymmetric sharing strategy to form the complete framework, Multi-$A$ Shared Adaptation (MASA). The overall architecture and data flow are illustrated in Fig. 4.
$h^{(l)}=W_{0}^{(l)}x+\frac{\alpha}{r}B^{(l)}\left(\sum_{i=1}^{N}A_{i}^{(k)}\right)x,\quad\text{where }k=\lfloor l/S\rfloor.$ (4)
Here, $k$ is the index of the shared $A$-matrix group, determined by the layer index $l$ (0-indexed) and the size of sharing group $S$.
Parameter Complexity Analysis: For a model with $L$ layers, the number of trainable parameters in MASA is approximately $\frac{L}{S}\times N\times d_{in}\times r+L\times d_{out}\times r$. The sharing mechanism significantly reduces the cost associated with the $A$-matrices, which would otherwise be the largest contributor due to multiple experts ($N$). This makes MASA’s parameter count comparable to or even lower than standard LoRA ($L\times(d_{in}+d_{out})\times r$), especially when $S$ is large. For example, on LLaMA3-8B, MASA in the $3A1B$ and $S=2$ configuration reduces $\sim$ 36.3% parameters than HydraLoRA in the corresponding $1A3B$ configuration.
Experiments
@@
Table 1: Comparison of zero-shot performance (%) of our proposed MASA method against other PEFT techniques on LLaMA3-8B. Performance is reported as accuracy on the MMLU benchmark and three single-domain tasks. The AA and $B$ specify the number of $A$ and $B$ matrices in the respective LoRA architectures. The best accuracy in each column is highlighted in bold. The results based on LLaMA3.2-3B are in Appendix D.
Experimental Setups
Datasets and Benchmarks. All experiments are conducted on the pretrained LLaMA3-8B, LLaMA3.1-8B and LLaMA3.2-3B models (Grattafiori et al. 2024). To evaluate multi-domain generalization, we fine-tune the model on the databricks-dolly-15k (Conover et al. 2023) dataset and assess its performance on the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al. 2020). For single-domain specialization, we focus on three distinct professional fields. In Law, the model is fine-tuned on Lawyer-Instruct (Alignment-Lab-AI 2024) and US-Terms (Li et al. 2023) and evaluated on the law-related tasks within MMLU. For Math, we use the training set of GSM8k (Cobbe et al. 2021) for fine-tuning and its test set for evaluation. Finally, in the Finance domain, we fine-tune on the training set of Fingpt-fineval (Wang, Yang, and Wang 2023) and test on its corresponding evaluation set. For multi-task reasoning, we fine-tune the model on a subset of the OpenOrca (Lian et al. 2023) dataset and evaluate its performance on the Big-Bench Hard (BBH) benchmark (Suzgun et al. 2022). Further details regarding these datasets are provided in Appendix C.
Implementation details. Unless otherwise noted, we fix the LoRA rank to $r=8$, scaling factor to $\alpha=16$, number of AA‑experts to $N=5$, and the size of sharing group to $S=2$. All models are fine‑tuned with AdamW ($\text{lr}=5^{-5}$, $\text{batch size}=8$). A full hyper‑parameter table is provided in Appendix C.
Baselines. To ensure a comprehensive evaluation, we compare the proposed MASA against representative methods spanning different PEFT paradigms:
  1. 1.
    Prompt and Representation Tuning: Methods that leave the backbone weights untouched and instead learn continuous prompts or prefix vectors: Prompt Tuning (Lester, Al-Rfou, and Constant 2021) and P-Tuning (Liu et al. 2021)
  2. 2.
    Single–Adapter LoRA Family: All methods insert one lightweight adapter per modified module, differing only in how the low‑rank update is parameterised: LoRA (Hu et al. 2022), DoRA (Liu et al. 2024), VB‑LoRA (Li, Han, and Ji 2024), and BS‑LoRA (Zhou et al. 2025)
  3. 3.
    Multi–Adapter and Modular LoRA: Approaches that instantiate multiple LoRA‑style adapters and decide at run‑time how to combine them: LoRAMoE (Dou et al. 2024), HydraLoRA (Tian et al. 2024), MTL‑LoRA (Yang et al. 2025), and CoLA (Zhou, Yao, and Chen 2025).
@@
Comparison results on Big-Bench Hard (BBH). To assess the model’s complex reasoning capabilities, we evaluated MASA on the Big-Bench Hard (BBH) benchmark, which is specifically designed to challenge multi-step reasoning abilities. As detailed in Tab. 2, MASA consistently achieves leading performance across different baselines, reaching a top accuracy of 42.82% on the LLaMA3.1-8B model while maintaining comparable learning parameters of $\sim$ 0.52%. Specifically, MASA demonstrates a clear advantage over alternative multi-expert designs that follow the “single-$A$, multi-BB” paradigm, such as HydraLoRA, and remains highly competitive with other methods like MTL-LoRA. The performance on BBH further validates that MASA’s “multi-$A$ architecture effectively alleviates the representational bottleneck, enhancing the model’s capacity for complex problem-solving.
Ablation Study and Analysis
@@
Model
MMLU Avg.
Params.(%)
Multi-AA
ACS.
@@
Analysis of optimal AA/$B$ configuration. We systematically explored the impact of different combinations of $A$ and $B$ matrices on the adaptation performance. As summarized in Tab. 4, we present the model’s accuracy on the MMLU benchmark as a function of various ($A$, $B$) configurations. Scaling $A$ with B=1B{=}1. Increasing $A$ from $3$ to $5$ improves MMLU performance to 59.62% with the modest parameter growth, while further scaling to A=6A{=}6 declines the performance, indicating the performance saturation. Scaling $B$ with A=5A{=}5. In contrast, increasing $B$ from $1$ to $3$ nearly doubles the parameter overhead while slightly reducing accuracy. This suggests that multiple $B$ projections introduce the redundancy to decrease performance. These results validate the “multi‑$A$, single‑BB” principle: a small ensemble of $A$ matrices can enrich the representation space, whereas duplicating $B$ mostly inflates the model.
Detailed effect of ASC. We further validate the rationale and optimality of our ASC. As the heatmaps in the methodology section revealed, $A$ matrices exhibit higher inter-layer similarity than $B$ matrices, which inspired our design to share only the $A$ matrices. To directly prove the superiority of this decision in terms of experimental performance, we conducted a comparative experiment across three different sharing strategies, which are summarized in Tab. 5.
The results confirm that our design is well-motivated. The “Share $A$, not BB” strategy achieves a favorable balance between performance and efficiency. In contrast, “Share $B$, not AA” leads to a significant performance drop, suggesting the importance of layer-specific $B$ matrices for model expressiveness. Despite sharing both $A$ and $B$ improving parameter efficiency, it degrades performance from 59.62% to 59.51%, indicating that excessive sharing of $B$ may hinder layer-specific transformations. Overall, Asymmetric Cross-layer Sharing on the $A$-matrices will be the best selection for adaptation.
@@ Conclusion
Appendix A Appendix
Appendix B Appendix A Proof of Theorem 1
We restate the theorem for convenience.
Theorem 2 (Information Ceiling).
Let $\mathbf{x}\!\in\!\mathbb{R}^{d_{\mathrm{in}}}$ be a zero-mean random vector with finite covariance $\Sigma_{\!x}\!=\!\mathbb{E}[\mathbf{x}\mathbf{x}^{\!\top}]\succeq 0$. Let the down-projection and the deterministic post-mapping be
$\mathbf{u}=A\mathbf{x},\quad\mathbf{v}=B\mathbf{u},\,A\!\in\!\mathbb{R}^{r\times d_{\mathrm{in}}},\;B\!\in\!\mathbb{R}^{d_{\mathrm{out}}\times r}.$
Then
$I(\mathbf{x};\mathbf{v})\;\leq\;\frac{1}{2}\log\!\det\!\bigl(I_{r}+A\Sigma_{\!x}A^{\!\top}\bigr),\,r(\mathbf{u})\leq r(A)\leq r.$} (5)
Proof.
Step 1 (Data-processing). Because $\mathbf{v}=B\mathbf{u}$ is a deterministic function of $\mathbf{u}$, the chain $\mathbf{x}\!\rightarrow\!\mathbf{u}\!\rightarrow\!\mathbf{v}$ is Markov; the data-processing inequality (DPI) gives:
$I(\mathbf{x};\mathbf{v})\;\leq\;I(\mathbf{x};\mathbf{u}).$ (6)
Step 2 (Gaussian upper envelope). Denote by $\mathbf{x}_{\mathrm{G}}\!\sim\!\mathcal{N}(0,\Sigma_{\!x})$ a Gaussian vector sharing the same second-order moment as $\mathbf{x}$, and set $\mathbf{u}_{\mathrm{G}}\!=\!A\mathbf{x}_{\mathrm{G}}$. The joint pair $(\mathbf{x}_{\mathrm{G}},\mathbf{u}_{\mathrm{G}})$ has the same mean, covariance and cross-covariance as $(\mathbf{x},\mathbf{u})$ because both involve the same linear map $A$. Under these matched second-order constraints, the Gaussian pair maximises mutual information, hence
$I(\mathbf{x};\mathbf{u})\;\leq\;I(\mathbf{x}_{\mathrm{G}};\mathbf{u}_{\mathrm{G}}).$ (7)
Step 3 (Closed form for linear–Gaussian). For jointly Gaussian variables linked by a deterministic linear map, $I(\mathbf{x}_{\mathrm{G}};\mathbf{u}_{\mathrm{G}})\!=\!\tfrac{1}{2}\log\det\!\bigl(I_{r}+A\Sigma_{\!x}A^{\!\top}\bigr)$ .
Step 4 (Chain the bounds). Combining (6)–(7) with Step 3 yields the desired mutual-information bound (5). The rank inequality follows from the sub-multiplicativity of rank: $r(\mathbf{u})=r(A\mathbf{x})\leq r(A)\leq r$. ∎
Remarks.
  1. 1.
    Tightness. Equality in (5) is attained when $(\mathbf{x},\mathbf{u})$ is itself jointly Gaussian, irrespective of $B$.
  2. 2.
    Why multi-$A$ lifts the ceiling. Replacing $A$ by an ensemble $\{A_{i}\}_{i=1}^{N}$ and aggregating $\tilde{\mathbf{u}}=\sum_{i}A_{i}\mathbf{x}$ preserves the rank constraint, but increases $\det\!\bigl(I_{r}+\!\sum_{i}A_{i}\Sigma_{\!x}A_{i}^{\!\top}\bigr)$ whenever the row-spaces of the $A_{i}$ are not perfectly aligned, thereby raising the right-hand side of (5).
Dataset Lang. Source / Version Task Types
[…]
Table 7: Overview of the datasets employed in this study.
Appendix C Appendix B Rationale for Summation-Based Aggregation

### v2

@@ 3.4 The Multi-$A$ Expert Block
As discussed above, our strategy is to reallocate the extra capacity to the feature extraction stage. The Multi-$A$ Expert (MAE) block replaces the single matrix $A$ with an ensemble of $N$ expert matrices, $\{A_{1},A_{2},\dots,A_{N}\}$, where each $A_{i}\in\mathbb{R}^{r\times d_{in}}$. These experts operate in parallel, enabling the model to capture more diverse and fine-grained features from the input. We hypothesize that each expert matrix learns to specialize in distinct semantic subspaces. This functional specialization is empirically supported by t-SNE (Maaten and Hinton, 2008) visualizations in Fig. 2, which illustrate that different $A_{i}$ matrices map the same inputs to separable feature clusters. The outputs of these experts are aggregated and then projected back to the original dimension by a single restoration matrix $B\in\mathbb{R}^{d_{out}\times r}$. We aggregate expert features via summation, a simple yet effective approach that maintains parameter efficiency and training stability (Xu et al., 2025), with detailed rationale provided in Appendix A.
@@
3.5 Asymmetric Cross-layer Sharing
To mitigate the potential parameter increase from using multiple $A$-matrices and to reduce their redundancy across transformer layers, we introduce an asymmetric cross-layer sharing mechanism for efficient adaptation. Unlike other sharing methods (VB-LoRA (Li et al., 2024), BSLoRA (Zhou et al., 2025b)) that typically share the symmetric LoRA module (both $A$ and $B$), our approach introduces an asymmetric sharing strategy: we only share the ensemble of $A$-matrices across adjacent layers, while allowing each layer to retain its own independent, trainable $B$-matrix.
To substantiate our design, we conducted an empirical analysis of inter-layer activation similarity using Centered Kernel Alignment (CKA) (Kornblith et al., 2019), a robust method for comparing neural network representations. As presented in Fig. 3, we applied CKA to measure the similarity between activation values from adjacent layers of a fine-tuned model, specifically comparing the $A$-matrix outputs ($Ax$) against the final LoRA increment activations ($BAx$). The analysis reveals a distinct pattern: The $A$-matrix outputs maintain a significantly higher similarity across layers than the full increments. This observation is quantified by the aggregated plots, which confirm a consistently higher average similarity score for the $A$-matrix outputs across all attention projections ($q,k,v$). This evidence indicates that 1) $A$-matrices learn to produce generalizable cross-layer representations, making them ideal candidates for sharing; and 2) $B$-matrices, conversely, perform more layer-specific transformations, which are critical for preserving the unique functionality of each layer.
@@ 3.5 Asymmetric Cross-layer Sharing
3.6 Discussion on the benefits of Multi-$A$ Shared Extractors
Consider a multi-branch extractor $\{A_{i}\}_{i=1}^{N}$ with summation aggregation:
$\sum_{i=1}^{N}A_{i}\mathbf{x}=A_{\text{sum}}\mathbf{x},\qquad A_{\text{sum}}\triangleq\sum_{i=1}^{N}A_{i}.$ (6)
Thus, under summation, multi-$A$ is functionally equivalent to a single extractor $A_{\text{sum}}$ and does not guarantee a larger ceiling in (3). Its benefit is trainability: splitting $A_{\text{sum}}$ into multiple branches changes regularization and optimization dynamics and promotes specialization under heterogeneous adaptation signals.
Weaker effective $\ell_{2}$ regularization on the realized extractor. With per-branch weight decay $\lambda\sum_{i=1}^{N}\|A_{i}\|_{F}^{2}$, the induced penalty on $A_{\text{sum}}$ is weaker: for any fixed $A_{\text{sum}}$, convexity implies
$\sum_{i=1}^{N}\|A_{i}\|_{F}^{2}\;\geq\;\frac{1}{N}\Big\|\sum_{i=1}^{N}A_{i}\Big\|_{F}^{2}\;=\;\frac{1}{N}\|A_{\text{sum}}\|_{F}^{2},$ (7)
with equality at $A_{i}=A_{\text{sum}}/N$. Hence multi-$A$ induces a weaker penalty on $A_{\text{sum}}$ in the optimal decomposition sense, allowing stronger extractor adaptation under the same rank-$r$ bottleneck.
Reduced destructive gradient interference in an expanded parameter space. In multi-task/domain adaptation, gradients w.r.t. a shared extractor can conflict. For a single $A$, the update aggregates task gradients $g=\sum_{t}g_{t}$, whose magnitude is diminished by negative cross terms: $\|g\|^{2}=\sum_{t}\|g_{t}\|^{2}+2\sum_{s<t}\langle g_{s},g_{t}\rangle$. With multiple branches, the over-parameterized decomposition provides additional degrees of freedom; under stochastic optimization, branches can follow different trajectories, which may alleviate cancellation in the effective update of the realized extractor.
The benefits of symmetry breaking.
With summation aggregation, each layer uses $A_{\text{sum}}^{(k)}=\sum_{i=1}^{N}A_{i}^{(k)}$ (Eq. 6), hence the objective is permutation-symmetric over experts. In practice, random initialization and stochastic optimization break this symmetry, and experts are not forced to become identical. For a layer $\ell$ in group $k$, $\Delta y_{\ell}=B_{\ell}A_{\text{sum}}^{(k)}x_{\ell}$ gives
$\nabla_{A_{i}^{(k)}}\mathcal{L}=\sum_{\ell\in\mathcal{G}(k)}\big(B_{\ell}^{\top}g_{\ell}\big)\,x_{\ell}^{\top},\qquad g_{\ell}\triangleq\nabla_{\Delta y_{\ell}}\mathcal{L},$ (8)
which is identical for all $i$ within the same group under summation. Nevertheless, the over-parameterized decomposition together with symmetry breaking can sustain expert diversity in practice. We empirically observe expert specialization (Fig. 2), which may help mitigate gradient interference under heterogeneous adaptation signals.
Low parameter complexity of MASA. For a model with $L$ layers, the number of trainable parameters in MASA is approximately $\frac{L}{S}\times N\times d_{in}\times r+L\times d_{out}\times r$. The sharing mechanism significantly reduces the cost associated with the $A$-matrices, which would otherwise be the largest contributor due to multiple experts ($N$). This makes MASA’s parameter count comparable to or even lower than standard LoRA ($L\times(d_{in}+d_{out})\times r$), especially when $S$ is large. For example, on LLaMA3-8B, MASA in the $3A1B$ and $S=2$ configuration reduces parameters by $\sim$ 36.3% compared to HydraLoRA in the corresponding $1A3B$ configuration.
4 Experiments
Table 1: Comparison of zero-shot performance (%) of our proposed MASA method against other PEFT techniques on LLaMA3-8B. Performance is reported as accuracy on the MMLU benchmark and three single-domain tasks. The AA and $B$ specify the number of $A$ and $B$ matrices in the respective LoRA architectures. The best accuracy in each column is highlighted in bold. The results based on LLaMA3.2-3B are in Appendix B.
@@ 4 Experiments
4.1 Experimental Setups
Datasets and Benchmarks. All experiments are conducted on the pretrained LLaMA3-8B, LLaMA3.1-8B and LLaMA3.2-3B models (Grattafiori et al., 2024). To evaluate multi-domain generalization, we fine-tune the model on the databricks-dolly-15k (Conover et al., 2023) dataset and assess its performance on the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2020). For single-domain specialization, we focus on three distinct professional fields. In Law, the model is fine-tuned on Lawyer-Instruct (Alignment-Lab-AI, 2024) and US-Terms (Li et al., 2023) and evaluated on the law-related tasks within MMLU. For Math, we use the training set of GSM8K (Cobbe et al., 2021) for fine-tuning and its test set for evaluation. Finally, in the Finance domain, we fine-tune on the training set of FinGPT-fineval (Wang et al., 2023a) and test on its corresponding evaluation set. For multi-task reasoning, we fine-tune the model on a subset of the OpenOrca (Lian et al., 2023) dataset and evaluate its performance on the Big-Bench Hard (BBH) benchmark (Suzgun et al., 2022). Further details regarding these datasets are provided in Appendix C.
Implementation details. Unless otherwise noted, we fix the LoRA rank to $r=8$, scaling factor to $\alpha=16$, number of AA‑experts to $N=5$, and the size of sharing group to $S=2$. All models are fine‑tuned with AdamW. A full hyper‑parameter table is provided in Appendix C.
Baselines. To ensure a comprehensive evaluation, we compare the proposed MASA against representative methods spanning different PEFT paradigms: (1) Prompt and Representation Tuning: Methods that leave the backbone weights untouched and instead learn continuous prompts or prefix vectors: Prompt Tuning (Lester et al., 2021) and P-Tuning (Liu et al., 2021). (2) Single–Adapter LoRA Family: All methods insert one lightweight adapter per modified module, differing only in how the low‑rank update is parameterised: LoRA (Hu et al., 2022), DoRA (Liu et al., 2024), VB‑LoRA (Li et al., 2024), and BS‑LoRA (Zhou et al., 2025b). (3) Multi–Adapter and Modular LoRA: Approaches that instantiate multiple LoRA‑style adapters and decide at run‑time how to combine them: LoRAMoE (Dou et al., 2024), HydraLoRA (Tian et al., 2024), MTL‑LoRA (Yang et al., 2025), and CoLA (Zhou et al., 2025a).
@@ 4.2 Main Results
Comparison results on Big-Bench Hard (BBH). To assess the model’s complex reasoning capabilities, we evaluated MASA on the Big-Bench Hard (BBH) benchmark, which is specifically designed to challenge multi-step reasoning abilities. As detailed in Tab. 2, MASA consistently achieves leading performance across different baselines, reaching a top accuracy of 42.82% on the LLaMA3.1-8B model while maintaining comparable parameter overhead of $\sim$ 0.52%. Specifically, MASA demonstrates a clear advantage over alternative multi-expert designs that follow the “single-$A$, multi-BB” paradigm, such as HydraLoRA, and remains highly competitive with other methods like MTL-LoRA. The performance on BBH further validates that MASA’s “multi-$A$ architecture effectively alleviates the representational bottleneck, enhancing the model’s capacity for complex problem-solving.
Additional LLM backbones. To further verify robustness beyond LLaMA3, we evaluate MASA on two additional instruction-tuned LLMs: Qwen2.5-7B-Instruct and Mistral-7B under the same MMLU protocol. On Qwen2.5-7B-Instruct, MASA outperforms HydraLoRA and LoRA by +0.44pp and +0.20pp in average MMLU, respectively. On Mistral-7B, MASA improves average MMLU by +0.81pp over HydraLoRA and +0.16pp over LoRA. Results on LLaMA3.2-3B show consistent trends, and full breakdowns are provided in Appendix B.
4.3 Ablation Study and Analysis
@@
Model MMLU Avg. Params. (%) Multi-AA ACS
@@
Analysis of optimal AA/$B$ configuration. We systematically explored the impact of different combinations of $A$ and $B$ matrices on the adaptation performance. As summarized in Tab. 4, we present the model’s accuracy on the MMLU benchmark as a function of various ($A$, $B$) configurations. Scaling $A$ with B=1B{=}1. Increasing $A$ from $3$ to $5$ improves MMLU performance to 59.62% with the modest parameter growth, while further scaling to A=6A{=}6 declines the performance, indicating the performance saturation. Scaling $B$ with A=5A{=}5. In contrast, increasing $B$ from $1$ to $3$ nearly doubles the parameter overhead while slightly reducing accuracy. This suggests that multiple $B$ projections introduce redundancy and slightly hurt performance. These results validate the “multi‑$A$, single‑BB” principle: a small ensemble of $A$ matrices can enrich the representation space, whereas duplicating $B$ mostly inflates the model.
Detailed effect of ACS. We further validate the rationale and optimality of our ACS. As the heatmaps in the methodology section revealed, $A$ matrices exhibit higher inter-layer similarity than $B$ matrices, which inspired our design to share only the $A$ matrices. To directly prove the superiority of this decision in terms of experimental performance, we conducted a comparative experiment across three different sharing strategies, which are summarized in Tab. 5.
The results confirm that our design is well-motivated. The “Share $A$, not BB” strategy achieves a favorable balance between performance and efficiency. In contrast, “Share $B$, not AA” leads to a significant performance drop, suggesting the importance of layer-specific $B$ matrices for model expressiveness. Despite sharing both $A$ and $B$ improving parameter efficiency, it degrades performance from 59.62% to 59.51%, indicating that excessive sharing of $B$ may hinder layer-specific transformations. Overall, Asymmetric Cross-layer Sharing on the $A$-matrices is the best selection for adaptation.
@@ 5 Conclusion
Appendix A Rationale for Summation-Based Aggregation